# Inferring neutron-star Love-Q relations from gravitational waves in the hierarchical Bayesian framework

### v1

@@ 1 Introduction
Future next-generation (XG) ground-based GW detectors, including the Cosmic Explorer (CE) [54, 55] and the Einstein Telescope (ET) [56, 57, 58, 59], are expected to detect many more GW signals, up to about $10^{5}$$10^{6}$ events per year [60, 61, 62, 63], thanks to their increased sensitivity and lower cutoff frequencies. These high-precision GW observations allow us to treat $\Lambda$ and $Q$ as independent parameters in the waveform model, to be measured directly from GWs. This enables further constraints on the Love-Q relation from an observational perspective. Samajdar and Dietrich [64] have first performed an analysis discussing the prospects of constraining Love-Q relation with GW observations, where a weighted linear regression was performed. While the results are very valuable, as we will show below, such a treatment might miss possible degeneracy and non-Gaussianity in the posteriors of $\Lambda$ and $Q$.
@@ 2.1 Polynomial Models of Love-Q Relation
Yagi and Yunes [38, 39, 43] fit the Love-Q relation with a quadratic polynomial model as,
@@ 2.2 Hierarchical Bayesian Inference
In the population inference scenario, the population properties are characterized by a set of hyperparameters, such as the power index of the mass distribution. These hyperparameters do not enter the waveform directly, while can be inferred from a collection of measurements of single-event parameters. Similarly, the EOS parameters determine the relation between $\Lambda$ and $m$, as well as between $Q$ and $m$, can be regarded as hyperparameters and analyzed in the hierarchical Bayesian framework [65, 66, 67, 68, 69, 70, 71, 72]. In this work, we adopt this framework to the inference of the Love-Q relation.
@@
Since we are inferring the Love-Q relation, the tidal deformabilities, $\bm{\Lambda}_{i}=\{\Lambda_{1i},\Lambda_{2i}\}$, and quadrupole moments, $\bm{Q}_{i}=\{Q_{1i},Q_{2i}\}$, of the two NSs of the $i\text{-th}$ event are treated as independent parameters in $\bm{\theta}_{i}$. Note that the Love-Q relation is insensitive to EOS, the priors of mass parameters, $m_{1}$ and $m_{2}$, are chosen to be independent of $\bm{\Lambda}_{i},\bm{Q}_{i}$ and $\bm{H}$. Other parameters in $\bm{\theta}_{i}$ are also assumed to be independent. In this way, the conditional prior of $\bm{\theta}_{i}$ can be further decomposed as
@@ 3.1 Waveform, Population and Detectors
In our simulation, we adopt the IMRPhenomXAS_NRTidalv3 waveform model [18], which includes tidal amplitude corrections as well as spin-induced quadrupole moment terms up to 3.5 PN with aligned spins. The single-event parameters $\bm{\theta}$ include the binary masses $m_{1}$ and $m_{2}$, the dimensionless tidal deformabilities $\Lambda_{1}$ and $\Lambda_{2}$, spin-induced quadrupole moments $Q_{1}$ and $Q_{2}$, the dimensionless spins $\chi_{1}$ and $\chi_{2}$, the luminosity distance $D_{L}$, the merge time $t_{c}$, the right ascension $\alpha$ and declination $\delta$, the inclination angle $\iota$, the GW polarization angle $\psi$, and the phase of coalescence $\phi_{c}$.
@@
We select a XG detector network consisting of two CE detectors and one ET detector, whose sensitivities are taken as CE-2 [54, 55] and ET-D [56, 57, 58], respectively. The two CE detectors are positioned at the current sites of the two LIGO detectors, while the ET detector is set at the current location of the Virgo detector with a triangular shape.
@@ 3.2 Implementation
We simulate 1000 GW events as described in section 3.1. Though the hierarchical procedure avoids a direct high-dimensional inference for $\bm{H}$ and $\{\bm{\theta}_{1},\cdots,\bm{\theta}_{n}\}$ simultaneously, the computational cost still increases with the number of events $n$. Lackey and Wade [68] found that when constraining the EOS of NSs, several events with the highest signal-to-noise ratios (SNRs) contribute most of the information. Also, Yagi and Yunes [39] concluded that the spin-induced quadrupole moment cannot be well measured for NSs with low spins. Considering these findings, we first draw 100 sources with the highest SNRs, then further select 20 sources with the largest $|\chi_{\mathrm{r}}|$ among them. As we will show in section 4, the primary information for constraining the Love-Q relation comes from the 10 loudest GW events. We leave a more comprehensive study of the whole population for future work.
@@ 4 Results and Discussions
In this section, we present the inference results for different parameterization models of the Love-Q relation. In section 4.1, we show the results of the linear model, Eq. (2.2), consisting of two parameters. In section 4.2, we present the results of the quartic polynomial model in Eq. (2.1), consisting of five parameters. The polynomial models in between, i.e., the quadratic and cubic polynomial models, are discussed in section 4.3.
@@ 4.1 The Linear Model
Figure 2: Posterior distributions of the hyperparameters ${\bm{H}}=\{a_{2},b_{2}\}$ in the linear fitting model. The contours refer to 50% and 90% credible regions, while the numbers above the histograms on the diagonal stand for the median and the central 90% credible interval of the marginalized distribution. We use blue and red colors to represent the results based on the loudest 20 and 10 events from the 1000 simulated events, respectively. The black lines represent the reference values in table 1. The grey lines on the diagonal represent the priors for comparison.
We first perform the inference for the linear model, which is also the model adopted in Ref. [64]. The posterior distribution of the hyperparameters $\{a_{2},b_{2}\}$ is shown in figure 2. Compared to the priors in table 1, the posteriors are significantly narrowed. As a comparison, we also mark the fitting values of the hyperparameters in a direct fit of the Yagi-Yunes Love-Q relation, and regard them as “true” values of the linear model for reference. These values fall within the 90% credible region of the posterior. The hierarchical Bayesian inference successfully recovers the Love-Q relation under the linear parameterization. We also investigate how the results depend on the number of events in the analysis. In the 20 events selected in section 3.2, we further select the loudest 10 events, perform the inference again, and show the results in figure 2. For both the joint and marginalized distributions, the widths of the credible regions in the 10-event inference are only slightly larger than those in the 20-event inference. This reveals that the loudest 10 events dominate the information in constraining the Love-Q relation, and including quieter events will not significantly change the results.
@@ 5 Testing Modified Gravity: Dynamical Chern-Simons Gravity
where $g$ is the determinant of the metric, $\kappa_{g}=1/16\pi$, $\mathcal{R}$ is the Ricci scalar, $\mathcal{R}_{\nu\mu\rho\sigma}$ and $^{*}\mathcal{R}^{\mu\nu\rho\sigma}$ denote the Riemann tensor and its dual, $\mathcal{L}_{\mathrm{mat}}$ is the matter Lagrangian density, $\alpha$ and $\beta$ are the coupling constants in the dCS gravity and the potential for the pseudo-scalar field $\mathcal{\vartheta}$ is omitted. $\mathcal{\vartheta}$ and $\beta$ is taken to be dimensionless, so $\alpha$ has the dimension of length square, and the quantity $\xi_{\mathrm{CS}}^{1/4}\equiv[\alpha^{2}/(\kappa\beta)]^{1/4}$ can be explained as the characteristic lengthscale of the dCS gravity [94, 95]. Current Solar System observations have constrained it to $\xi_{\mathrm{CS}}^{1/4}<\mathcal{O}(10^{8})$ km [96, 95].
@@ 6 Conclusion
There can be several extensions to this study. Firstly, we assume BNS systems with aligned spins, while realistic BNS systems may have tilted spins and spin precession, which leads to additional contribution in the GW waveform [99, 100, 101]. Additionally, in the XG era, the GW signals are possibly overlapping with each other, complicating the data analysis [102, 63, 103, 104, 105]. Moreover, the Yagi-Yunes universal relations were originally found and discussed for a single NS. For BNS systems, these universal relations may hold when the two NSs have similar masses, but still require a careful inspection [1, 106]. The octupole or higher-order multipole moments of NSs also contribute to the GW waveform and can be universally related to the quadrupole moment [43, 18]. We leave the exploration of simultaneous or independent inference of more universal relations to future studies.
@@ Acknowledgments
This work was supported by the Beijing Natural Science Foundation (QY25102, 1242018), the National Natural Science Foundation of China (123B2043, 12573042), the National SKA Program of China (2020SKA0120300), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-Performance Computing Platform of Peking University.

### v2

@@ 1 Introduction
Future next-generation (XG) ground-based GW detectors, including the Cosmic Explorer (CE) [76, 77] and the Einstein Telescope (ET) [74, 46, 82, 3], are expected to detect many more GW signals, up to about $10^{5}$$10^{6}$ events per year for BNS coalescences [8, 83, 52, 81], thanks to their increased sensitivity and lower cutoff frequencies. These high-precision GW observations allow us to treat $\Lambda$ and $Q$ as independent parameters in the waveform model, to be measured directly from GWs. This enables further constraints on the Love-Q relation from an observational perspective. Samajdar and Dietrich [80] have first performed an analysis discussing the prospects of constraining Love-Q relation with GW observations, where a weighted linear regression was performed. While the results are very valuable, as we will show below, such a treatment might miss possible degeneracy and non-Gaussianity in the posteriors of $\Lambda$ and $Q$.
@@ 2.1 Polynomial Models of Love-Q Relation
Yagi and Yunes [108, 107, 110] fit the Love-Q relation with a quartic polynomial model as,
@@ 2.2 Hierarchical Bayesian Inference
In the population inference scenario, the population properties are characterized by a set of hyperparameters, such as the power index of the mass distribution. These hyperparameters do not enter the waveform directly, while can be inferred from a collection of measurements of single-event parameters. The EOS parameters, which determine the $\Lambda$-$m$ and $Q$-$m$ relations, can also be regarded as hyperparameters and analyzed in the hierarchical Bayesian framework [62, 63, 10, 56, 61, 40, 9, 95]. In this work, we adopt this framework to the inference of the Love-Q relation.
@@
We treat the tidal deformabilities, $\bm{\Lambda}_{i}=\{\Lambda_{1i},\Lambda_{2i}\}$, and quadrupole moments, $\bm{Q}_{i}=\{Q_{1i},Q_{2i}\}$, of the two NSs of the $i\text{-th}$ event as independent parameters in $\bm{\theta}_{i}$ and select flat priors for them. The impacts of priors are relatively small, since only the loudest events with high SNRs are analyzed. Note that the Love-Q relation is insensitive to EOS, the priors of mass parameters, $m_{1}$ and $m_{2}$, are chosen to be independent of $\bm{\Lambda}_{i},\bm{Q}_{i}$ and $\bm{H}$. Other parameters in $\bm{\theta}_{i}$ are also assumed to be independent. In this way, the conditional prior of $\bm{\theta}_{i}$ can be further decomposed as
@@ 3.1 Waveform, Population and Detectors
In our simulation, we adopt the IMRPhenomXAS_NRTidalv3 waveform model [2], which includes tidal amplitude corrections as well as spin-induced quadrupole moment terms up to 3.5 PN with aligned spins. The single-event parameters $\bm{\theta}$ include the binary masses $m_{1}$ and $m_{2}$, the dimensionless tidal deformabilities $\Lambda_{1}$ and $\Lambda_{2}$, spin-induced quadrupole moments $Q_{1}$ and $Q_{2}$, the dimensionless spins $\chi_{1}$ and $\chi_{2}$, the luminosity distance $D_{L}$, the coalescence time $t_{c}$, the right ascension $\alpha$ and declination $\delta$, the inclination angle $\iota$, the GW polarization angle $\psi$, and the phase of coalescence $\phi_{c}$.
@@
where $\alpha=0.68$, $\mu_{1}=1.34\,\mathrm{M}_{\odot}$, $\sigma_{1}=0.02\,\mathrm{M}_{\odot}$, $\mu_{2}=1.47\,\mathrm{M}_{\odot}$, $\sigma_{2}=0.15\,\mathrm{M}_{\odot}$, and $m_{\mathrm{s}}^{l}=1.14\,\mathrm{M}_{\odot}$, $m_{\mathrm{s}}^{u}=1.46\,\mathrm{M}_{\odot}$. The NS with a larger mass is labeled as the primary star with mass $m_{1}$ and the other as the secondary star with mass $m_{2}$. The tidal deformability and quadrupole moments of the binary are calculated from the stellar mass assuming the APR4 EOS with methods described in Refs. [107, 19] and the slow rotation approximation is used. For arbitary spin, Refs. [29, 71, 103, 28] have developed new relations insensitive to certain set of EOSs with fitting coefficients depending on the spin parameter. Ref. [22] further suggested introducing the fitting coefficients as a function of both spin and radius, extending the universality to various EOSs. In these cases, our framework will still be useful with appropriate model parameterizations and modifications.
@@
We select a XG detector network consisting of two CE detectors and one ET detector, whose sensitivities are taken as CE-2 [76, 77] and ET-D [74, 46, 82], respectively. The two CE detectors are positioned at the current sites of the two LIGO detectors, while the ET detector is set at the current location of the Virgo detector with a triangular shape. The current locations of LIGO and Virgo are chosen as an illustrated configuration for XG detectors.
@@ 3.2 Implementation
We simulate 1000 GW events as described in section 3.1. Though the hierarchical procedure avoids a direct high-dimensional inference for $\bm{H}$ and $\{\bm{\theta}_{1},\cdots,\bm{\theta}_{n}\}$ simultaneously, the computational cost still increases with the number of events $n$. Lackey and Wade [56] found that when constraining the EOS of NSs, several events with the highest signal-to-noise ratios (SNRs) contribute most of the information. Also, Yagi and Yunes [107] concluded that the spin-induced quadrupole moment cannot be well measured for NSs with low spins. Considering these findings, we first draw 100 sources with the highest SNRs, then further select 20 sources with the largest $|\chi_{\mathrm{r}}|$ among them. The SNR values of these selected events range from about 1500 to 2500. Here we omit the selection effects since different $\Lambda$ and $Q$ values do not lead to a significant change of SNR and thus the detection probability. As we will show in section 4, the primary information for constraining the Love-Q relation comes from the loudest GW events. In practice, selecting the top 20 events is enough to capture the main information for constraining the Love-Q relation. We present the results of changing the number of selected events in appendix A. We leave a more comprehensive study of the whole population considering selection effects for future work.
@@ 4 Results and Discussions
In this section, we present the inference results for different parameterization models of the Love-Q relation. In section 4.1, we show the results of the linear model, Eq. (2.2), consisting of two parameters. In section 4.2, we present the results of the quartic polynomial model in Eq. (2.1), consisting of five parameters. The polynomial models in between, i.e., the quadratic and cubic polynomial models, are discussed in section 4.3. For other parameterizations beyond polynomial models, our methodology is in principle applicable to them as well.
@@ 4.1 The Linear Model
Figure 2: Posterior distributions of the hyperparameters ${\bm{H}}=\{a_{2},b_{2}\}$ in the linear fitting model. The contours refer to 50% and 90% credible regions, while the numbers above the histograms on the diagonal stand for the median and the central 90% credible interval of the marginalized distribution. We use blue and red colors to represent the results based on the loudest 20 and 10 events from the 1000 simulated events, respectively. The orange bands represent the reference values with fitting uncertainties accounted for. The grey lines on the diagonal represent the priors for comparison.
We first perform the inference for the linear model, which is also the model adopted in Ref. [80]. The posterior distribution of the hyperparameters $\{a_{2},b_{2}\}$ is shown in figure 2. Compared to the priors in table 1, the posteriors are significantly narrowed. As a comparison, we also mark the fitting values of the hyperparameters in a direct fit of the Yagi-Yunes Love-Q relation, and regard them as “true” values of the linear model for reference. These values fall within the 90% credible region of the posterior. Given that the Love-Q relation is quasi-universal with relative differences of $\sim 1\%$ for various EOSs, the fitting uncertainties of the hyperparameters are also considered in the reference values. To check the robustness of our conclusions, we also repeat the inference assuming the SLy EOS; which are summarized in appendix B. The hierarchical Bayesian inference successfully recovers the Love-Q relation under the linear parameterization. We also investigate how the results depend on the number of events in the analysis. In the 20 events selected in section 3.2, we further select the loudest 10 events, perform the inference again, and show the results in figure 2. For both the joint and marginalized distributions, the widths of the credible regions in the 10-event inference are only slightly larger than those in the 20-event inference. This reveals that the loudest 10 events dominate the information in constraining the Love-Q relation, and including quieter events will not significantly change the results.
@@ 5 Testing Modified Gravity: Dynamical Chern-Simons Gravity
where $g$ is the determinant of the metric, $\kappa_{g}=1/16\pi$, $\mathcal{R}$ is the Ricci scalar, $\mathcal{R}_{\nu\mu\rho\sigma}$ and $^{*}\mathcal{R}^{\mu\nu\rho\sigma}$ denote the Riemann tensor and its dual, $\mathcal{L}_{\mathrm{mat}}$ is the matter Lagrangian density, $\alpha$ and $\beta$ are the coupling constants in the dCS gravity and the potential for the pseudo-scalar field $\mathcal{\vartheta}$ is omitted. $\mathcal{\vartheta}$ and $\beta$ are taken to be dimensionless, so $\alpha$ has the dimension of length squared, and the quantity $\xi_{\mathrm{CS}}^{1/4}\equiv[\alpha^{2}/(\kappa\beta)]^{1/4}$ can be explained as the characteristic lengthscale of the dCS gravity [112, 106]. Current Solar System observations have constrained it to $\xi_{\mathrm{CS}}^{1/4}<\mathcal{O}(10^{8})$ km [16, 106].
@@ 6 Conclusion
There can be several extensions to this study. Firstly, we assume BNS systems with aligned spins, while realistic BNS systems may have tilted spins and spin precession, which leads to additional contribution in the GW waveform [102, 75, 37]. The inaccuracy of waveform itself can also introduce a theoretical error to the posteriors of single event parameters, especially for events with high SNRs [25]. Additionally, in the XG era, the GW signals are possibly overlapping with each other, complicating the data analysis [72, 81, 97, 51, 96]. Moreover, the Yagi-Yunes universal relations were originally found and discussed for a single NS under the slow-rotation assumption. For NSs in BNS systems or with high spins, these universal relations may break down and new relations could emerge. Our methodology in principle could still work when new universal relations do exist, but a careful inspection is required [85, 78, 29, 71, 22, 103, 28, 101]. The octupole or higher-order multipole moments of NSs also contribute to the GW waveform and can be universally related to the quadrupole moment [110, 2]. We leave the exploration of simultaneous or independent inference of more universal relations to future studies.
@@ Acknowledgments
We thank the anonymous referee for helpful comments. This work was supported by the Beijing Natural Science Foundation (QY25102, 1242018), the National Natural Science Foundation of China (123B2043, 12573042), the National SKA Program of China (2020SKA0120300), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-Performance Computing Platform of Peking University.
Appendix A Dependence on the Number of Events
In the main text, we have selected the loudest 20 events from the 1000 simulated events in the inference. Here we vary the number of events $N_{\text{events}}$ to see how the results change. As shown in figure 7, for a small $N_{\text{events}}$ (like $N=5$), there is a noticeable offset between the median estimate and the reference value of the hyperparameters. However, as $N_{\text{events}}$ increases, both the median and the credible intervals shrink towards the reference values. to the reference values. Due to the decreasing SNRs of the latter added samples, their contribution to the inference becomes less and less significant, and the results keep almost unchanged. The choice of 20 loudest events can be justified with figure 8, where we observe that the constraint to Love-Q relation at $\Lambda\sim 350$ (the “waist” part in figure 3 of the manuscript) becomes stable when $N_{\text{events}}$ is larger than 20. Similar results are also found in previous studies of constraining the EOSs with GWs, where the several loudest events dominate the information [56].
Figure 7: The dependency of the 5th, 50th and 95th percentiles of the hyperparameter posterior samples on the number of events. The orange dashed line represents the reference values, i.e. the linear fitting values obtained in section 2.1.
Figure 8: The dependency of 5th and 95th percentiles of $Q$ samples computed at $\Lambda=350$ (where most of the data points gather) on the number of events. The orange dashed line represents $Q$ calculated from the Yagi-Yunes relation.
Appendix B Results for SLy
Different assumptions of EOS provide different combinations of $\Lambda$ and $Q$, which still follow the universal relation. As a supplement, we have repeated the inference of linear model for SLy, another EOS consistent with GW170817 [4, 5, 6]. The results are presented and compared with those of APR4 in figure 9 and figure 10.
As shown in figure 9, the uncertainties of fitting coefficients brought by multiple EOSs are much smaller compared to those from the hierarchical inference. Thus the change of EOS would not introduce a significant change to the width of constraints, while the median may probably change with EOSs. And in figure 10. The median and maximum posterior estimations change, but the width of 90% credible region remains insensitive to the choice of EOS.
Figure 9: Corner plots for APR4 and SLy EOSs.
Figure 10: Constraints to Love-Q relation in a linear model for APR4 and SLy.